OpenStudy (anonymous):

If f(x) = 3x + 1, then f(a + h) - f(a) =

OpenStudy (anonymous):

f(a+h)=3(a+h)+1=3a+3h+1 f(a)=3a+1 now f(a+h)-f(a)=3a+3h+1-(3a+1) =3a+3h+1-3a-1=3h answer.